# Neural Force Field: Few-shot Learning of Generalized Physical Reasoning

### v4

@@ Abstract
Physical reasoning is a remarkable human ability that enables rapid learning and generalization from limited experience. Current AI models, despite extensive training, still struggle to achieve similar generalization, especially in Out-of-distribution (OOD) settings. This limitation stems from their inability to abstract core physical principles from observations. A key challenge is developing representations that can efficiently learn and generalize physical dynamics from minimal data. Here we present Neural Force Field (NFF), a framework extending Neural Ordinary Differential Equation (NODE) to learn complex object interactions through force field representations, which can be efficiently integrated through an Ordinary Differential Equation (ODE) solver to predict object trajectories. Unlike existing approaches that rely on discrete latent spaces, NFF captures fundamental physical concepts such as gravity, support, and collision in continuous explicit force fields. Experiments on three challenging physical reasoning tasks demonstrate that NFF, trained with only a few examples, achieves strong generalization to unseen scenarios. This physics-grounded representation enables efficient forward-backward planning and rapid adaptation through interactive refinement. Our work suggests that incorporating physics-inspired representations into learning systems can help bridge the gap between artificial and human physical reasoning capabilities.
@@ 1 Introduction
Physical reasoning, the ability to understand and predict how objects interact in the physical world, is fundamental to both human intelligence and artificial systems (Spelke, , 2022). This capability underlies crucial applications ranging from robotics to scientific discovery, making it a central challenge in AI research (Lake et al., , 2017). One of the remarkable aspects of human cognitive capabilities is the ability to rapidly learn from limited examples (Kim et al., , 2020; Jiang et al., , 2022; Lake and Baroni, , 2023; Zhang et al., , 2024), especially evident in intuitive physics (Kubricht et al., , 2017; Bear et al., , 2021). Humans can quickly abstract core physical principles after observing limited physical phenomena, enabling them to predict complex dynamics and interact with novel environments (Spelke and Kinzler, , 2007; Battaglia et al., , 2013; Bonawitz et al., , 2019; Xu et al., , 2021).
In contrast, current AI systems face significant limitations in physical reasoning. Despite being trained on gigantic datasets, these models still struggle to achieve human-level generalization, particularly in Out-of-distribution (OOD) settings (Lake et al., , 2017; Zhu et al., , 2020). The core issue lies in their tendency to overfit observed trajectories rather than capturing inherent physical principles, severely limiting their ability to compose known knowledge and predict outcomes in novel contexts (Qi et al., , 2021; Li et al., , 2022; Wu et al., , 2023). This stark contrast between human capabilities and current model limitations has motivated the search for new approaches that can learn generalizable physical representations from minimal data.
@@
To address these challenges, we introduce Neural Force Field (NFF), a neural function map from objects to forces acting on continuous Euclidean coordinates for efficient interactive learning and reasoning. At its core, NFF employs a neural operator to learn dynamic force fields from external interventions and object interactions. These predicted forces are then integrated through an ODE solver to compute explicit physical variables such as velocity and displacement, producing interpretable results that align with established physical principles. Please refer to Fig.˜1 for details.
@@
We evaluate our method using two state-based physical reasoning benchmarks—I-PHYRE (Li et al., , 2023) and N-body problems (Newton, , 1833)—as well as one vision-based benchmark, PHYRE (Bakhtin et al., , 2019). These tasks feature complex rigid-body dynamics ranging from short-range forces (collision and sliding) to long-range forces (gravity). Our experiments demonstrate that NFF not only learns dynamics efficiently by abstracting physical interactions into force fields but also achieves strong generalization in both within-scenario and cross-scenario settings. Moreover, the framework’s physics-based representation enables effective forward and backward planning in goal-directed tasks, consistently outperforming existing GNN-based methods such as Interaction Network (IN) (Battaglia et al., , 2016; Qi et al., , 2021) and transformer-based methods (Wu et al., , 2023).
@@ 2 Related work
Physical reasoning  Research in physical reasoning has progressed along two main trajectories: passive observation and interactive platforms. The passive observation approach, exemplified by the Violation-of-Expectation (VoE) paradigm (Spelke et al., , 1992), evaluates physical understanding by measuring agents’ ability to detect violations of physics principles (Baillargeon, , 1987; Hespos and Baillargeon, , 2001; Dai et al., , 2023). While this approach has provided valuable insights into basic physical comprehension, it is limited in assessing active interventions and complex reasoning.
@@
To enable more comprehensive evaluation, interactive platforms such as PHYRE (Bakhtin et al., , 2019), the virtual tools game (Allen et al., , 2020), and I-PHYRE (Li et al., , 2023) have emerged. These environments require agents to actively manipulate objects to achieve specific goals, testing not only prediction capabilities but also planning and reasoning skills. However, existing approaches in these platforms often struggle with generalization across diverse scenarios, particularly in few-shot settings where limited training data is available (Qi et al., , 2021; Li et al., , 2023).
Current physical reasoning models face two primary challenges: the need for extensive training data and limited cross-scenario transferability. While some methods achieve strong performance within specific scenarios (Allen et al., , 2020), they often fail to generalize their understanding to novel situations, falling short of human-level reasoning capabilities (Kang et al., , 2025). Our work addresses these limitations by introducing a framework that unifies passive observation and interactive learning through dynamic force fields. The NFF approach enables few-shot learning of physical principles while supporting active reasoning through its interpretable force-based representation, facilitating both accurate prediction and effective intervention planning across diverse physical scenarios.
Dynamic prediction  The prediction of physical dynamics, fundamental to intuitive physics, has evolved along two methodological branches: discrete and continuous-time approaches. Discrete methods typically combine recurrent architectures with GNNs (Battaglia et al., , 2016; Qi et al., , 2021) or transformer modules (Wu et al., , 2023) for modeling object interactions, while leveraging convolutional operators for processing pixel-based information (Shi et al., , 2015; Wang et al., , 2022). Despite their flexibility in handling various interaction types, these methods often struggle with two key limitations: difficulty in extracting robust physical representations and susceptibility to error accumulation over extended time horizons.
Continuous-time methods address the above issues by incorporating physical inductive biases via explicit modeling of state derivatives within dynamical systems (Chen et al., , 2018; Zhong et al., , 2020; Norcliffe et al., , 2020). While they improve long-term prediction, current systems typically assume energy-conservative systems with simplified dynamics (Greydanus et al., , 2019; Cranmer et al., , 2020) or static fields (Kofinas et al., , 2023). More importantly, these methods face significant challenges in few-shot learning and struggle with cross-scenario generalization (Chen et al., , 2020; Yuan et al., , 2024), often requiring specific priors and grammars to achieve meaningful transfer (Xu et al., , 2021). Although Physics-informed Neural Networks (PINN) can be precise in modeling physical dynamics, they depend heavily on prior knowledge of the underlying equations (Chu et al., , 2022).
@@
Our work advances continuous-time methods to model complex rigid body interactions via learnable dynamic force fields. This approach enables robust cross-scenario generalization while handling non-conservative energy systems. By combining the stability advantages of continuous-time modeling with flexible force field representations, NFF achieves both accurate long-term predictions and strong generalization capabilities without requiring explicit physical priors.
@@ 3.1 The Neural Force Field (NFF) representation
Traditional approaches to modeling physical interactions typically rely on implicit representations through hidden vectors to describe object interactions. These methods use neural networks to learn state transition functions that map current scene features to future states. However, this purely data-driven approach cannot guarantee physically grounded representations, leading to poor generalization despite intensive training. We introduce NFF, which addresses these limitations by learning physics-grounded, generalizable representations through explicit force field modeling; see Fig.˜2 for a comparison.
Predicting force field  The core of NFF is built on the physical concept of a force field—a vector field that determines the force experienced by any query object based on its state $\mathbf{z}^{q}$. For a scene with $N$ objects, we model the force field $\mathbf{F}(\cdot)$ at time $t$ using a neural network operating on a relation graph $\mathcal{G}$. The graph contains object nodes $V=\{\mathbf{z}^{0}(t),\mathbf{z}^{1}(t),...,\mathbf{z}^{N-1}(t)\}$. Following neural operator methods (Lu et al., , 2021) and graph neural models (Battaglia et al., , 2018), we formulate the force field function as:
$\small\mathbf{F}(\mathbf{z}^{q}(t))=\sum_{i\in\mathcal{G}(q)}\mathbf{W}\left(f_{\theta}(\mathbf{z}^{i}(t))\odot f_{\phi}(\mathbf{z}^{q}(t))\right)+\mathbf{b},$ (1)
where $\mathcal{G}(q)$ represents the neighbors of query $q$ determined by contacts or attractions, $f_{\theta}$ and $f_{\phi}$ are neural networks with parameters $\theta$ and $\phi$, and $\mathbf{W}\in\mathbb{R}^{d_{\text{hidden}}\times d_{\text{force}}},\mathbf{b}\in\mathbb{R}^{d_{\text{force}}}$ map hidden features to low-dimensional forces. The state vector $\mathbf{z}$ incorporates zero-order (position, angle) and first-order (velocity, angular velocity) states. This representation can handle various physical interactions including collision, sliding, rotation, and gravity, as shown in Fig.˜1.
Figure 2: Comparison between discrete latent transition in traditional interaction modeling and continuous explicit transition in NFF. Traditional interaction models such as IN and SlotFormer decode discrete frames from a learnable decoder. In contrast, NFF decodes trajectories by integrating the continuous force using an ODE solver, which benefits learning detailed interactions such as collisions.
@@
$\small\frac{d^{2}x^{q}(t)}{dt^{2}}=\frac{dv^{q}(t)}{dt}\propto\mathbf{F}(\mathbf{z}^{q}(t)).$ (2)
While object mass theoretically affects acceleration, these mass constants are safely absorbed into the learned $F(\cdot)$ either by being explicitly included as inputs or implicitly captured through correlated physical attributes like size.
@@
$\small\mathbf{z}^{q}(t)=\text{ODESolve}\left(\mathbf{z}^{q}(0),\mathbf{F},0,t\right),$ (3)
$\small\begin{dcases}\mathbf{x}(t)=\mathbf{x}(0)+\int_{0}^{t}\mathbf{v}(t)\,dt,\\ \mathbf{v}(t)=\mathbf{v}(0)+\int_{0}^{t}\mathbf{F}(\mathbf{z}^{q}(t))\,dt,\end{dcases}$ (4)
The NFF framework offers three key advantages over existing approaches. First, it enables few-shot learning of force fields from minimal examples—even single trajectories—through its low-dimensional, physically-grounded representation $\mathbf{F}(\mathbf{z}^{q}(t))$, adhering to Eq.˜2. Second, by representing interactions as neural operator force fields, as shown in Eq.˜1, NFF can generalize the learned force patterns to novel scenarios through simple force summation on new interaction graphs, allowing modeling varying object compositions, interaction types, and time horizons. Third, the high-precision integration decoder, formulated in Eqs.˜3 and 4, enables both fine-grained modeling of continuous-time interactions and efficient optimization for planning tasks through its invertible information flow. This precision supports various applications, from determining initial conditions for desired outcomes to designing targeted force fields for specific behaviors. The effectiveness of these three components—ODE grounding, Neural Operator Learning (NOL), and precise integration—is systematically validated through ablation studies in Sec.˜4.5.
@@ 3.2 Interactive planning
Forward planning  For physical reasoning tasks requiring specific action sequences, we extend the NFF framework to incorporate action effects by including action features in $f_{\theta}$. Through forward simulation with the NFF model, we sample multiple action sequences $A=\{a_{0},...,a_{T}\}$ and optimize according to evaluation metrics $R$: $A^{\star}=\underset{A}{\text{argmax}}R(\mathbf{F},A)$. The optimal sequence $A^{\star}$ can then be executed in the physical environment. While the forward simulation may initially deviate from actual observations, NFF’s physics-based design enables efficient adaptation through new experimental data, inspired by human trial-and-error learning (Allen et al., , 2020). When executed trajectories deviate from the desired goal state, the new state sequences can be directly incorporated into model optimization through the MSE loss, following the same training procedure used in the initial learning.
Backward planning  The invertible nature of the NFF formulation makes backward computation particularly efficient. By inverting the time direction of ODE integrator, NFF enables the inversion of initial conditions given a desired goal state: $\mathbf{x}(0)=\mathbf{x}(t)+\int_{t}^{0}\mathbf{v}(t)\,dt$, $\mathbf{v}(0)=\mathbf{v}(t)+\int_{t}^{0}\mathbf{F}(\mathbf{z}^{q}(t))\,dt$, consistent with the learned physical dynamics encoded in the NFF model, as shown in Fig.˜10.
4 Experiments
4.1 Environments and datasets
We evaluate state-based NFF on I-PHYRE and N-body dynamics and vision-based NFF on PHYRE. Each task is evaluated across three distinct settings: within-scenario prediction, cross-scenario prediction, and planning. Detailed environment and dataset configurations are provided in Appx.˜A.
I-PHYRE   I-PHYRE (Li et al., , 2023) presents a suite of complex physical reasoning puzzles requiring multi-step interventions. The environment incorporates diverse physical interactions including gravity, collision, friction, rotation, spring dynamics, and pendulum motion. It challenges AI agents to solve puzzles with minimal environmental interactions while generalizing to unseen scenarios. For within-scenario prediction, we evaluate on 10 training games that share similar scenarios but require different solutions. The cross-scenario prediction setting extends to 30 novel games featuring noise, compositional elements, and multi-ball scenarios, with varying object properties such as block lengths, ball sizes, and object positions. In the planning setting, models must generate optimal action sequences to successfully complete each game.
N-body  The N-body problem (Newton, , 1833) tests trajectory prediction for small comets orbiting a massive central planet. Using REBOUND simulator (Rein and Liu, , 2012), we generate dynamics by randomly sampling orbital parameters including radii, angles, and masses. This task evaluates the model’s ability to infer gravitational laws from limited observations. The within-scenario prediction introduces novel initial conditions and masses, focusing on systems with 1-2 orbiting bodies tracked over 50 timesteps. Cross-scenario prediction significantly increases complexity by introducing systems with 7 or 9 orbiting bodies tracked over 150 timesteps. The planning setting challenges models to optimize initial conditions that will evolve to specified target states after 50 timesteps.
PHYRE  Building on the previous two datasets that used state-based representations of objects, we explore extending NFF to a vision-based physical reasoning benchmark, PHYRE (Bakhtin et al., , 2019). Designed for goal-oriented physical reasoning, PHYRE features 25 task templates involving interactive physics puzzles, all solvable by strategically placing a red ball. Each template generates 100 variations of a core task, supporting two distinct evaluation protocols: (i) the within-scenario setting, where models train on 80% of the tasks in each template and are tested on the remaining 20%; and (ii) the cross-scenario setting, which evaluates generalization by training on all tasks from 20 templates and testing on tasks from the 5 unseen ones.

### v5

@@ Abstract
Physical reasoning is a remarkable human ability that enables rapid learning and generalization from limited experience. Current AI models, despite extensive training, still struggle to achieve similar generalization, especially in Out-of-distribution (OOD) settings. This limitation stems from their inability to abstract core physical principles from observations. A key challenge is developing representations that can efficiently learn and generalize physical dynamics from minimal data. Here we present Neural Force Field (NFF), a framework extending Neural Ordinary Differential Equation (NODE) to learn complex object interactions through force field representations, which can be efficiently integrated through an Ordinary Differential Equation (ODE) solver to predict object trajectories. Unlike existing approaches that rely on discrete latent spaces, NFF captures fundamental physical concepts such as gravity, support, and collision in continuous explicit force fields. Experiments on three challenging abstract physical reasoning tasks demonstrate that NFF, trained with only a few examples, achieves strong generalization to unseen scenarios. This physics-grounded representation enables efficient forward-backward planning and rapid adaptation through interactive refinement. Our work suggests that incorporating physics-inspired representations into learning systems can help bridge the gap between artificial and human physical reasoning capabilities.
@@ 1 Introduction
Physical reasoning, the ability to understand and predict how objects interact in the physical world, is fundamental to both human intelligence and artificial systems (Spelke, 2022). This capability underlies crucial applications ranging from robotics to scientific discovery, making it a central challenge in AI research (Lake et al., 2017). One of the remarkable aspects of human cognitive capabilities is the ability to rapidly learn from limited examples (Kim et al., 2020; Jiang et al., 2022; Lake & Baroni, 2023; Zhang et al., 2024), especially evident in intuitive physics (Kubricht et al., 2017; Bear et al., 2021). Humans can quickly abstract core physical principles after observing limited physical phenomena, enabling them to predict complex dynamics and interact with novel environments (Spelke & Kinzler, 2007; Battaglia et al., 2013; Bonawitz et al., 2019; Xu et al., 2021).
In contrast, current AI systems face significant limitations in physical reasoning. Despite being trained on gigantic datasets, these models still struggle to achieve human-level generalization, particularly in Out-of-distribution (OOD) settings (Lake et al., 2017; Zhu et al., 2020). The core issue lies in their tendency to overfit observed trajectories rather than capturing inherent physical principles, severely limiting their ability to compose known knowledge and predict outcomes in novel contexts (Qi et al., 2021; Li et al., 2022; Wu et al., 2023). This stark contrast between human capabilities and current model limitations has motivated the search for new approaches that can learn generalizable physical representations from minimal data.
@@
To address these challenges, we introduce Neural Force Field (NFF), a neural function map from objects to forces acting on continuous Euclidean coordinates for efficient interactive learning and reasoning. At its core, NFF employs a neural operator to learn dynamic force fields from external interventions and object interactions. These predicted forces are then integrated through an ODE solver to compute explicit physical variables such as velocity and displacement, producing interpretable results that align with established physical principles. Please refer to Figure 1 for details.
@@
We evaluate our method using two state-based physical reasoning benchmarks—I-PHYRE (Li et al., 2023) and N-body problems (Newton, 1833)—as well as one vision-based benchmark, PHYRE (Bakhtin et al., 2019). These tasks feature complex rigid-body dynamics ranging from short-range forces (collision and sliding) to long-range forces (gravity). Our experiments demonstrate that NFF not only learns dynamics efficiently by abstracting physical interactions into force fields but also achieves strong generalization in both within-scenario and cross-scenario settings. Moreover, the framework’s physics-based representation enables effective forward and backward planning in goal-directed tasks, consistently outperforming existing methods such as RPIN (Qi et al., 2021), EGNN (Satorras et al., 2021), SEGNO (Liu et al., 2024b), and transformer-based methods (Wu et al., 2023).
@@ 2 Related work
Physical reasoning   Research in physical reasoning has progressed along two main trajectories: passive observation and interactive platforms. The passive observation approach, exemplified by the Violation-of-Expectation (VoE) paradigm (Spelke et al., 1992), evaluates physical understanding by measuring agents’ ability to detect violations of physics principles (Baillargeon, 1987; Hespos & Baillargeon, 2001; Dai et al., 2023). While this approach has provided valuable insights into basic physical comprehension, it is limited in assessing active interventions and complex reasoning.
@@
To enable more comprehensive evaluation, interactive platforms such as PHYRE (Bakhtin et al., 2019), the virtual tools game (Allen et al., 2020), and I-PHYRE (Li et al., 2023) have emerged. These environments require agents to actively manipulate objects to achieve specific goals, testing not only prediction capabilities but also planning and reasoning skills. However, current physical reasoning models face two primary challenges: the need for extensive training data (Qi et al., 2021; Li et al., 2023) and limited cross-scenario transferability. While some methods achieve strong performance within specific scenarios (Allen et al., 2020), they often fail to generalize their understanding to novel situations, falling short of human-level reasoning capabilities (Kang et al., 2025).
Object-centric Dynamic prediction   The prediction of physical dynamics between objects is fundamental to intuitive physics. Many approaches (Battaglia et al., 2016; Qi et al., 2021) adopt Graph Neural Networks (GNN) as their backbone, capitalizing on their relational inductive bias for modeling object-centric interactions. Key enhancements include integrating E(3) equivariant architectures to generalize across rotations and translations (Satorras et al., 2021; Han et al., 2022), and incorporating NODE for intrinsic continuity modeling to improve long-term prediction accuracy (Poli et al., 2019; Yuan et al., 2024; Liu et al., 2024b). Some works inject physics inductive bias into simulation such as mesh or SDF (Allen et al., 2023; Rubanova et al., 2024) for collision handling and spring-mass models or particle-grid representations (Jiang et al., 2025; Zhang et al., 2025) for deformable objects. However, these methods are mainly designed for particle-based systems which hardly fit into abstract physical reasoning tasks like PHYRE. More importantly, these methods fail to extract robust physical representations, which severely hinders their performance in few-shot learning scenarios. More recent methods like SlotFormer (Wu et al., 2023) and SlotSSM (Jiang et al., 2024) leverage transformer or mamba for temporal sequence modeling with a slot attention (Locatello et al., 2020) module for unsupervised object detection, but still struggle with cross-scenario generalization.
@@
Our work advances continuous-time methods to model complex rigid body interactions via learnable dynamic force fields. This approach enables robust cross-scenario generalization while handling non-conservative energy systems. By combining the stability advantages of continuous-time modeling with flexible force field representations, NFF achieves both accurate long-term predictions and strong generalization capabilities without requiring explicit physical priors. More related work can be found in Appendix A.
@@ 3.1 The Neural Force Field (NFF) representation
Traditional approaches to modeling physical interactions typically rely on implicit representations through hidden vectors to describe object interactions. These methods use neural networks to learn state transition functions that map current scene features to future states. However, this purely data-driven approach cannot guarantee physically grounded representations, leading to poor generalization despite intensive training. We introduce NFF, which addresses these limitations by learning physics-grounded, generalizable representations through explicit force field modeling; see Figure 2 for a comparison.
Predicting force field   The core of NFF is built on the physical concept of a force field—a vector field that determines the force experienced by any query object based on its state $\mathbf{z}^{q}$. For a scene with $N$ objects, we model the force field $\mathbf{F}(\cdot)$ at time $t$ using a neural network operating on a relation graph $\mathcal{G}$. The graph contains object nodes $V=\{\mathbf{z}^{0}(t),\mathbf{z}^{1}(t),...,\mathbf{z}^{N-1}(t)\}$. Following neural operator methods (Lu et al., 2021) and graph neural models (Battaglia et al., 2018), we formulate the force field function as:
$\mathbf{F}(\mathbf{z}^{q}(t))=\sum_{i\in\mathcal{G}(q)}\mathbf{W}\left(f_{\theta}(\mathbf{z}^{i}(t))\odot f_{\phi}(\mathbf{z}^{q}(t))\right)+\mathbf{b},$ (1)
where $\mathcal{G}(q)$ represents the neighbors of query $q$ determined by contacts or attractions, $f_{\theta}$ and $f_{\phi}$ are neural networks with parameters $\theta$ and $\phi$, and $\mathbf{W}\in\mathbb{R}^{d_{\text{hidden}}\times d_{\text{force}}},\mathbf{b}\in\mathbb{R}^{d_{\text{force}}}$ map hidden features to low-dimensional forces. The state vector $\mathbf{z}$ incorporates geometry states (angle, length, radius), optionally physical attributes (mass), zero-order states (position), and first-order states (velocity, angular velocity). This representation can handle various physical interactions including collision, sliding, rotation, and gravity, as shown in Figure 1. For image-based NFF, we utilize object masks to replace the geometry states in input.
@@
$\mathbf{a}^{q}(t)=\frac{d^{2}x^{q}(t)}{dt^{2}}=\frac{dv^{q}(t)}{dt}=\frac{\mathbf{F}(\mathbf{z}^{q}(t))}{m^{q}}.$ (2)
While Equation 2 explicitly incorporates object mass $m^{q}$, our formulation is flexible regarding mass availability. In scenarios with known physical properties, $m^{q}$ serves as an explicit divisor. Conversely, when mass is not provided as a state input, the network $\mathbf{F}(\cdot)$ learns to implicitly encode mass information into its parameters via end-to-end optimization. In this implicit setting, $\mathbf{F}(\cdot)$ effectively predicts mass-normalized forces (i.e., accelerations) by leveraging correlated features such as object geometry or semantics.
Figure 2: Comparison between discrete latent transition in traditional interaction modeling and continuous explicit transition in NFF. The figures in the first line compare the predicted discrete steps and continuous trajectories, in which the discrete decoding cannot explain how the green ball goes through the black wall. In the second line, the black arrows indicate velocity vectors and the red and green arrows indicate force vectors. Traditional interaction models such as IN and SlotFormer decode discrete frames from a learnable decoder. In contrast, NFF decodes trajectories by integrating the continuous force using an ODE solver, which benefits learning detailed interactions such as collisions.
@@
$\mathbf{z}^{q}(t)=\text{ODESolve}\left(\mathbf{z}^{q}(0),\mathbf{F},0,t\right),$ (3)
$\begin{dcases}\mathbf{x}(t)=\mathbf{x}(0)+\int_{0}^{t}\mathbf{v}(t)\,dt,\\ \mathbf{v}(t)=\mathbf{v}(0)+\int_{0}^{t}\frac{\mathbf{F}(\mathbf{z}^{q}(t))}{m^{q}}\,dt,\end{dcases}$ (4)
The NFF framework offers three key advantages over existing approaches. First, it enables few-shot learning of force fields from minimal examples—even single trajectories—through its low-dimensional, physically-grounded representation $\mathbf{F}(\mathbf{z}^{q}(t))$, adhering to Equation 2. Second, by representing interactions as neural operator force fields, as shown in Equation 1, NFF can generalize the learned force patterns to novel scenarios through simple force summation on new interaction graphs, allowing modeling varying object compositions, interaction types, and time horizons. Third, the high-precision integration decoder, formulated in Equations 3 and 4, enables both fine-grained modeling of continuous-time interactions and efficient optimization for planning tasks through its invertible information flow. This precision supports various applications, from determining initial conditions for desired outcomes to designing targeted force fields for specific behaviors. The effectiveness of these three components—ODE grounding, Neural Operator Learning (NOL), and precise integration—is systematically validated through ablation studies in Section 4.5.
@@ 3.2 Interactive planning
Forward planning   For physical reasoning tasks requiring specific action sequences, we extend the NFF framework to incorporate action effects by including action features in $f_{\theta}$. Through forward simulation with the NFF model, we sample multiple action sequences $A=\{a_{0},...,a_{T}\}$ and optimize according to evaluation metrics $R$: $A^{\star}=\underset{A}{\text{argmax}}R(\mathbf{F},A)$. The optimal sequence $A^{\star}$ can then be executed in the physical environment. While the forward simulation may initially deviate from actual observations, NFF’s physics-based design enables efficient adaptation through new experimental data, inspired by human trial-and-error learning (Allen et al., 2020). When executed trajectories deviate from the desired goal state, the new state sequences can be directly incorporated into model optimization through the MSE loss, following the same training procedure used in the initial learning.
Backward planning   The invertible nature of the NFF formulation makes backward computation particularly efficient. By inverting the time direction of ODE integrator, NFF enables the inversion of initial conditions given a desired goal state: $\mathbf{x}(0)=\mathbf{x}(t)+\int_{t}^{0}\mathbf{v}(t)\,dt$, $\mathbf{v}(0)=\mathbf{v}(t)+\int_{t}^{0}\mathbf{F}(\mathbf{z}^{q}(t))\,dt$, consistent with the learned physical dynamics encoded in the NFF model, as shown in Figure 10.
@@
4 Experiments
4.1 Environments and datasets
We evaluate state-based NFF on I-PHYRE and N-body dynamics and vision-based NFF on PHYRE. Each task is evaluated across three distinct settings: within-scenario prediction, cross-scenario prediction, and planning. Detailed environment and dataset configurations are provided in Appendix B.
I-PHYRE    I-PHYRE (Li et al., 2023) presents a suite of complex physical reasoning puzzles requiring multi-step interventions. The environment incorporates diverse physical interactions including gravity, collision, friction, rotation, spring dynamics, and pendulum motion. It challenges AI agents to solve puzzles with minimal environmental interactions while generalizing to unseen scenarios. For within-scenario prediction, we evaluate on 10 training games that share similar scenarios but require different solutions. The cross-scenario prediction setting extends to 30 novel games featuring noise, compositional elements, and multi-ball scenarios, with varying object properties such as block lengths, ball sizes, and object positions. In the planning setting, models must generate optimal action sequences to successfully complete each game.
N-body   The N-body problem (Newton, 1833) tests trajectory prediction for small comets orbiting a massive central planet. Using REBOUND simulator (Rein & Liu, 2012), we generate dynamics by randomly sampling orbital parameters including radii, angles, and masses. This task evaluates the model’s ability to infer gravitational laws from limited observations. The within-scenario prediction introduces novel initial conditions and masses, focusing on systems with 1-2 orbiting bodies tracked over 50 timesteps. Cross-scenario prediction significantly increases complexity by introducing systems with 7 or 9 orbiting bodies tracked over 150 timesteps. The planning setting challenges models to optimize initial conditions that will evolve to specified target states after 50 timesteps.
PHYRE   Building on the previous two datasets that used state-based representations of objects, we explore extending NFF to a vision-based physical reasoning benchmark, PHYRE (Bakhtin et al., 2019). Designed for goal-oriented physical reasoning, PHYRE features 25 task templates involving interactive physics puzzles, all solvable by strategically placing a red ball. Each template generates 100 variations of a core task, supporting two distinct evaluation protocols: (i) the within-scenario setting, where models train on 80% of the tasks in each template and are tested on the remaining 20%; and (ii) the cross-scenario setting, which evaluates generalization by training on all tasks from 20 templates and testing on tasks from the 5 unseen ones.